# Rethinking Machine Unlearning for Large Language Models

### v3

@@ 6 Applications of LLM Unlearning
In this section, we discuss two application categories facilitated by LLM unlearning: the first focused on data influence and the second on model capabilities.
Copyright and privacy protection.
@@
Similar to deleting copyrighted information from the training data, another scenario is preventing LLMs from leaking user privacy, especially personal identification information (PII). This concern is closely related to LLM memorization and training data extraction (Carlini et al., 2019, 2021, 2022; Jang et al., 2022; Nasr et al., 2023).
Sociotechnical harm reduction.
Another application of LLM unlearning is alignment (Ouyang et al., 2022), aimed at aligning LLMs with human instructions and making sure generated text conforms to human values. Unlearning can be used to forget harmful behaviors such as the production of toxic, discriminatory, illegal, or morally undesirable outputs (Shevlane et al., 2023), e.g., instructions to build CBRN (chemical, biological, radiological, and nuclear) weapons. Unlearning, as a safety alignment tool, can happen at the different stages of LLM development, e.g., before, during, or after alignment. Current research has focused on the ‘pre-alignment’ stage (Yao et al., 2023), there may be untapped opportunities in the others.
@@
LLMs are known to generate biased decisions and outputs (Perez et al., 2022; Tamkin et al., 2023; Cui et al., 2023). In the vision domain, unlearning has proven to be an effective tool for reducing discrimination to enable fair decision-making (Sattigeri et al., 2022; Chen et al., 2023b). In the language domain, unlearning has been applied to mitigate gender-profession bias (Yu et al., 2023) and many other fairness issues (Sattigeri et al., 2022; Oesterling et al., 2023; Kadhe et al., 2023). However, more opportunities exist, such as unlearning stereotypes in training data.
LLMs are also known to be vulnerable to jailbreaking attacks (Wei et al., 2023; Qi et al., 2023; Huang et al., 2023) (i.e., adversarially crafted prompts that lead the LLM to generate undesired outputs) as well as poisoning/backdoor attacks (Rando & Tramèr, 2023; Carlini et al., 2023). Unlearning can be a natural solution for both types of attacks given the existing success of unlearning as a defense against adversarial attacks in other domains (Wang et al., 2019; Li et al., 2021; Liu et al., 2022; Jia et al., 2023).
7 Discussion and Conclusion
This position paper rethinks the paradigm of unlearning for modern LLMs to uncover its under-explored aspects. To achieve this, we dissect LLM unlearning into four essential aspects: formulation, methodologies, evaluation metrics, and applications. We show that there are considerable challenges in both foundational research and practical, use case-driven research. These include: (Generality) A desired solution for LLM unlearning should take into account the generality of the unlearning target and dataset choice, accommodate various model setups including both white-box and black-box scenarios, and consider the specifics of the unlearning method. (Authenticity) LLM unlearning should focus on effectively removing both data influence and specific model capabilities, in order to authenticate unlearning across a range of evaluation methods, particularly in adversarial contexts. (Precision) LLM unlearning should precisely define the scope of unlearning, while ensuring the preservation of general language modeling performance outside this unlearning scope.
@@ 7 Discussion and Conclusion
8 Broader Impacts
The broader impact of our work, particularly its ethical implications and societal consequences, revolves around the responsible use of data and maintaining the integrity of large language models (LLMs) and other foundational AI models. Our research aims to explore machine unlearning methods to enhance safety and security, uphold user privacy, reduce biases, and guarantee the generation of trustworthy information by AI systems. The methodologies and frameworks we explored have the potential to significantly influence ethical AI practices in the future. We recognize, however, that the societal implications of progress in machine unlearning are intricate and diverse. This necessitates ongoing dialogue and scrutiny, especially in relation to the dynamic field of AI ethics and governance. As machine unlearning finds application in practical scenarios, it is essential to thoughtfully address any ethical dilemmas and societal ramifications that may arise. This underscores our dedication to advancing AI technology in a manner that is both responsible and mindful of its broader impacts.

### v4

@@ 6 Applications of LLM Unlearning
There mainly exist two application areas (EAs) facilitated by LLM unlearning: the first focused on data influence and the second on model capabilities.
EA1: Copyright and privacy protection
@@
Similar to deleting copyrighted information from the training data, another scenario is preventing LLMs from leaking user privacy, especially personal identification information (PII). This concern is closely related to LLM memorization and training data extraction (Carlini et al., 2019, 2021, 2022; Jang et al., 2022; Nasr et al., 2023). Fig. 2 (row 1) provides concrete examples of LLM’s outputs post unlearning focusing on the removal of data influence for PII protection (Wu et al., 2023b) and copyrighted information degeneration (Eldan & Russinovich, 2023).
Figure 2: Demonstration of responses of LLMs post-unlearning across different applications. (Row 1) LLM unlearning focuses on removing undesired data influence. Left: PII (personal identification information) removal using the unlearning method PO (reject-based preference optimization) (Maini et al., 2024). Right: Degenerating copyrighted information in the Harry Potter book series (Eldan & Russinovich, 2023) using the unlearning method GradDiff (gradient difference of forget and retain losses) (Liu et al., 2022a; Maini et al., 2024). (Row 2) LLM unlearning focuses on model behavior alignment. Left: Prevention of generating toxic content using the unlearning method PO. Right: Prevention of the malicious use of LLMs, using the unlearning method NPO (negative preference optimization) (Zhang et al., 2024b), for generating hazardous knowledge.
EA2: Sociotechnical harm reduction
Another application of LLM unlearning is alignment (Ouyang et al., 2022), aimed at aligning LLMs with human instructions and making sure generated text conforms to human values. Unlearning can be used to forget harmful behaviors such as the production of toxic, discriminatory, illegal, or morally undesirable outputs (Shevlane et al., 2023; Gehman et al., 2020; Li et al., 2024a), e.g., instructions to build CBRN (chemical, biological, radiological, and nuclear) weapons. Unlearning, as a safety alignment tool, can happen at the different stages of LLM development, e.g., before, during, or after alignment. Current research has focused on the ‘pre-alignment’ stage (Yao et al., 2023), there may be untapped opportunities in the others. Fig. 2 (row 2) exemplifies the response of unlearned LLMs in detoxification (Yao et al., 2023) and reducing malicious use of LLMs on the WMDP benchmark(Li et al., 2024a).
@@
LLMs are known to generate biased decisions and outputs (Perez et al., 2022; Tamkin et al., 2023; Cui et al., 2023). In the vision domain, unlearning has proven to be an effective tool for reducing discrimination to enable fair decision-making (He et al., 2019; Sattigeri et al., 2022; Chen et al., 2023b; Dreyer et al., 2024). In the language domain, unlearning has been applied to mitigate gender-profession bias (Yu et al., 2023) and many other fairness issues (Sattigeri et al., 2022; Oesterling et al., 2023; Kadhe et al., 2023). However, more opportunities exist, such as unlearning stereotypes in training data.
LLMs are also known to be vulnerable to jailbreaking attacks (Wei et al., 2023; Qi et al., 2023; Huang et al., 2023; Zou et al., 2023) (i.e., adversarially crafted prompts that lead the LLM to generate undesired outputs) as well as poisoning/backdoor attacks (Rando & Tramèr, 2023; Carlini et al., 2023; Hubinger et al., 2024). Unlearning can be a natural solution for both types of attacks given the existing success of unlearning as a defense against adversarial attacks in other domains (Wang et al., 2019; Li et al., 2021; Liu et al., 2022b; Jia et al., 2023).
7 Challenges and Overlook
This work rethinks the paradigm of unlearning for modern LLMs to uncover its under-explored aspects. To achieve this, we dissect LLM unlearning into four essential aspects: formulation, methodologies, evaluation metrics, and applications. We show that there are considerable challenges in both foundational research and practical, use case-driven research. These include: (Generality) A desired solution for LLM unlearning should take into account the generality of the unlearning target and dataset choice, accommodate various model setups including both white-box and black-box scenarios, and consider the specifics of the unlearning method. (Authenticity) LLM unlearning should focus on effectively removing both data influence and specific model capabilities, in order to authenticate unlearning across a range of evaluation methods, particularly in adversarial contexts. (Precision) LLM unlearning should precisely define the scope of unlearning, while ensuring the preservation of general language modeling performance outside this unlearning scope.
@@ 7 Challenges and Overlook
In addition, the necessity for regulations or policies to govern unlearning practices is crucial for the future, given the potential implications on privacy, security, and fairness. While existing research has primarily concentrated on auditing unlearning processes related to membership inference, addressing this issue presents an immensely complex challenge. It includes a multitude of factors, including but not limited to data handling/attribution, model governance, transparency, accountability, and verification throughout the unlearning lifecycle. Regulations and policies need to address issues such as data retention, consent management, and the right to be forgotten, which are particularly critical in sensitive domains like healthcare and security applications.
Another future effort we encourage is to build an ‘LLM Unlearning Algorithm Card’ that carefully details the involved parties, data aimed to be unlearned, evaluation reports, and the implementation details of the unlearning practice.